OpenStudy (anonymous):

-3sin(theta)+3=2cos^2(theta)

OpenStudy (campbell_st):

well replace cos^2(theta) with 1-sin^2(theta) and you get a quadratic equation \[-3\sin(\theta) + 3 = 2(1 - \sin^2(\theta))\] or \[2\sin^2(\theta) - 3\sin(\theta) + 1 = 0\] this factors to \[(2\sin(\theta) - 1)(\sin(\theta) -1) = 0\] now you can solve the equation

OpenStudy (campbell_st):

hope it helps